OpenStudy (anonymous):

A solution contains 72.0 g of HCl and 468 g of (C6H6). What is the mole fraction of benzene?

OpenStudy (abmon98):

Molar Mass of HCL:(35.5+1)=36.5 grams Molar Mass of C6H6:(12*6)+(1*6)=78 grams Number of Moles:Mass(g)/Molar Mass(g/mol) 72/36.5=1.97260 moles of HCL 468/78=6moles of C6H6 Sum of Moles of Solution=6+1.97260 Moles of Solution=7.97260 Mole Fraction of benzene: Number of moles of Benzene/Number of Moles of Solution 6/7.97260

OpenStudy (anonymous):